Grade 12Physical Chemistry

4 gram of divalent metal was dissolved in 100 c.c of 0.1 N h2so4 the excess of acid required 30 c.c. of 1 N NaOH for complete neutralization. Find the atomic mass of metal ?

himal
8 Years agoGrade 12

1 Answer

Jenny Ynnej
4 Years ago
Soln:
Volume of Naoh(V1)= 30cc
Normality of Naoh(N1)= 1N
Volume of excess H2so4(V2) = ??
Normality of H2so4(N2) = molarity × basicity
= 2×2 = 4 × 1.01 = 4.04N
Using,
V1N1 = V2N2
30×1 = V2 × 4.04
I.e. V2 = 7.42cc 
Volume of excess acid = 7.42 cc
   Volume of H2so4 neutralized by metal=    100 - 7.42
= 92.575 cc
Therefore,  gm.eqv of H2so4 = gm.eqv of metal 
(92.575/1000) × 4.04 = 4/E
I.e E = 10.09 
Atomic mass= E × basicity
                       = 10.09 × 2
                       =20.18amu ans